OpenStudy (thatkid):

3x+4=5x-7 explain step by step please

OpenStudy (anonymous):

1) subtract 3x from both sides 2) add 7 to both sides 3) divide both sides by 2

OpenStudy (anonymous):

3x+4=5x-7 subtract 3x from both sides 4=2x-7 add 7 to both sides 11=2x divide by 2 x= 5.5